# Is Limited Information Enough? An Approximate Multi-agent Coverage Control in Non-Convex Discrete Environments

### 08 Jan 2024 v1

@@ Abstract.
Conventional distributed approaches to coverage control may suffer from lack of convergence and poor performance, due to the fact that agents have limited information, especially in non-convex discrete environments. To address this issue, we extend the approach of Marden (2016) which demonstrates how a limited degree of inter-agent communication can be exploited to overcome such pitfalls in one-dimensional discrete environments. The focus of this paper is on extending such results to general dimensional settings. We show that the extension is convergent and keeps the approximation ratio of 2, meaning that any stable solution is guaranteed to have a performance within 50% of the optimal one. We also show that the computational complexity and communication complexity are both polynomial in the size of the problem. The experimental results exhibit that our algorithm outperforms several state-of-the-art algorithms, and also that the runtime is scalable as per theory.
@@ 1. Introduction
It is known that, even in a centralized context, finding an optimal solution for the coverage problem is an NP-hard problem Megiddo and Supowit (1984). Hence, most studies focus on approximate approaches. In distributed settings, game-theoretical control approaches seek to design agents that will be incentivized to behave autonomously in a way that is well-aligned with the designer’s objective. This strategy has proven to be successful in a number of applications (see, e.g. Douchan et al. (2019)). The situation is even more difficult in practice since agents often have restricted sensing and communication capabilities. Consequently, agents must make decisions based on local information about their environment and the other agents. Unfortunately, algorithms based on local information may suffer from lack of convergence and degraded performance due to miscoordination between the agents. As for the convergence issue, it is known that a move of an agent can affect the cost of agents outside of the neighborhood, and thus, the decrease of the global cost cannot be guaranteed locally, especially in the case of discrete non-convex environment Yun and Rus (2014). The degradation of performance can be explained with a worst-case scenario in which only a single agent can perceive a large number of valuable locations within an environment, while a number of other agents cannot perceive these locations.
@@
Indeed, different settings exist that vary according to the information available:
@@
    agents may not communicate any information and be only guided by their local perception;
@@
    agents may communicate candidate individual moves to their neighbors only;
@@
    agents may communicate candidate moves (possibly coordinated with other agents) to agents (possibly beyond their neighbors).
@@
Existing game-theoretical approaches can be classified into these types: classical Voronoi-based best-response approaches fall into either (1) or (2), depending on the assumption. Two recent approaches in type (3) explore different directions. Sadeghi et al. Sadeghi et al. (2020) proposed a distributed algorithm for non-convex discrete settings in which agents have the possibility to coordinate a move with a single other agent (meaning that an agent moves and assumes another agent takes her place simultaneously), possibly beyond their neighbors, when individual moves are not sufficient. However, the algorithm sacrifices the approximation ratio for convergence. Marden Marden (2016) on the other hand allows coordinated moves with several agents, but only under the restriction that these agents are within the neighborhood. While this approach achieves a convergent algorithm with an approximation ratio of 2, by sharing only the information of the minimum utility among agents, it is limited in that the only investigated case is the one-dimensional (line) environment. This severely limits real-world applications. Other studies try to achieve global optimality by developing approaches akin to simulated annealing. For example, Arslan et al. (2007) attains global optimum with Spatial Adaptive Play (SAP, a.k.a Boltzmann exploration) and uses random search to escape from the local optimum. However, this approach suffers from a slow convergence rate when the search space is large. There is also no discussion about the relationship between information and efficiency. Note that we focus on the coverage problem, and multiagent path-planning issues Galceran and Carreras (2013); Stern et al. (2019); Okumura and Défago (2022) are out of the scope of the paper.
@@
In this paper, we extend the algorithm of Marden (2016) to any-dimensional, non-convex discrete space and compare this approach with the aforementioned alternative variants of game-theoretical control. Our approach can also be considered as a generalization of Durham et al. (2011), while they are limited to pairwise neighborhood without a guarantee of approximation ratio. The remainder of this paper is as follows. Section 2 introduces the model and existing approaches. In Section 3 we detail the algorithm and prove that it guarantees convergence to a neighborhood optimum solution with an approximation ratio of 2 without any restriction on the dimensionality of the environment, i.e., the same guarantee as in the 1D case. Additionally, we propose a polynomial-time extension of the algorithm and discuss the computation and communication complexity. Experiments reported in Section 4 indeed show that our algorithm outperforms existing ones, along with adhering to the theoretical approximation ratio. Furthermore, the runtime results confirm the scalability of the proposed algorithm.
@@ 2.1. Coverage Problems
We start with a set of agents $\mathcal{N}=\{1,\ldots,n\}$ and a set of resources $C=\{c_{1},\ldots,c_{m}\}$. In a coverage problem, resources are locations (or points) in a connected metric space. We assume that this is discrete finite space that is modeled as a connected graph $(C,\mathcal{E})$, where $\mathcal{E}$ is the set of edges that connect two adjacent points. Though our approach can be extended to continuous settings, we omit the detail due to the space limit. We denote the distance between two points $a,b\in C$ as $|a-b|$ that is the length of the shortest path.
@@
Let $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ denote a non-increasing function, $v_{c}\in\mathbb{R}_{+}$ be the weight of point $c\in C$, and $\mathcal{C}\subseteq C$ be a partial space. Then, the objective function for $\mathcal{C}$ is defined as follows:
@@
Let us then consider the coverage problem depicted in Figure 1. For the sake of exposure, let us assume that $g(d)=1/(1+d)$, and ||$\cdot$|| ⋅ | is Manhattan distance. Agents are identified by letters $a,b,\dots f$. The environment is a grid world, where circled locations are valued $v_{c}=1$, while the others are valued $v_{c}=0$. The locations covered by an agent are represented in grey and we indicate the name of the agent. For unoccupied locations we indicate on the node the (Manhattan) distance to the closest agent. Thus a covered location gives utility of 1, while a location at a distance of 1 from the closest agent gives utility 1/2, and so on. The depicted allocation $x$ has value $G(x)=1+1+\frac{1}{2}+\dots=16.4$. It is clearly sub-optimal. The reader can check that the allocation $x^{\prime}$ where agent $e$ has moved one location up would induce $G(x^{\prime})\simeq 16.8$.
@@ 2.2. Game-Theoretic Control: Generalities
This paper will focus on providing each agent $i$ with a utility function of the form $u_{i}:\mathcal{X}\to\mathbb{R}$ that will ultimately guide their individual behavior. In the rest of the paper, we formulate the set of choices of agent $i$ as $\mathcal{X}_{i}(x_{-i})$ assuming that it depends on a given allocation of the other agents $x_{-i}$. For simplicity, we denote $\mathcal{X}_{i}(x)=\mathcal{X}_{i}(x_{-i})$. We also denote $\mathcal{X}(x)=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\times}\mathcal{X}_{i}(x)$.
@@
In general, the efficiency of a Nash equilibrium $G(x)$ can be smaller than the optimal value $G(x^{*})$. One of the reasons of this suboptimality is the miscoordination among self-interested agents, as agents are required to make independent decisions in response to available information. Other sources of suboptimality comes from the structure of the agents’ utility functions and available choice sets. The (worst-case, among all instances) ratio between the worst Nash equilibria and the social optimum is known as the price of anarchy (PoA) Koutsoupias and Papadimitriou (1999). Here, the choice set $\mathcal{X}_{i}$ could encode physical constraints on choices, i.e. an agent can only physically choose a given subset of choices given the behavior of the collective $x$. Alternatively, the choice set could encode information availability of the agent, e.g. it is the set of choices for which the agent can evaluate its utility. Regardless of the interpretation, it is important to highlight that the structure of the choice sets can significantly alter the structure and efficiency of the resulting equilibria which we discuss in the ensuing section.
@@ 2.3. Local Information
Several approaches in game theory seek to exploit the fact that agents typically have limited sensing power and only a local view of the situation. Marden Marden (2016) analyzed how miscoordination among agents with limited information leads to inefficient Nash equilibria. To this end, he introduced the concept of information set which is the set of choices each agent can perceive and compute the resulting utilities. In this paper, $\mathcal{X}_{i}(x)$ corresponds to the information set that is the set of locations agent $i$ can perceive based on spatial proximity. With this notion, an allocation is a Nash equilibrium (Equation (3)) if agents are at least as happy as any choice for which they can evaluate their utility. Observe that the information set $\mathcal{X}_{i}(x)$ is a state-dependent notion: the local information available may vary depending on the current allocation $x$.
@@ 2.4. Linking Information and Inefficiency
Intuitively, local information can cause distributed systems based on game-theoretic control to get stuck in an inefficient allocation. Indeed, some agents who may obtain higher utilities for a resource may not have access to this information. The redundancy index hence gives insights about the amount of information available to the agents and guarantees on the worst-case ratio.
Marden Marden (2016) investigated this interplay, in the broader context of resource allocation games. Assuming that the global objective function $G$ is monotone submodular i.e., satisfies the following conditions:
@@
The following theorem shows how the value of $f$ impacts the efficiency of Nash equilibrium allocation:
@@
Thus, we know that the approximation ratio of a distributed allocation algorithm can be 2 in the case of full information ($f=n$), because $G(x)\geq\frac{1}{2}G(x^{*})$ in this case (from Equation (8)). Also, it is impossible to guarantee an approximation ratio better than $n$ in case of $f=1$ (from Equation (9)). We will use this result in the later section to analyze the efficiency of coverage control algorithms.
@@ 2.5. Voronoi-Based Control
One of the most standard approaches of coverage control is the algorithms based on Voronoi partitioning Gao et al. (2008). A Voronoi partition divides the space into local regions for each agent. Formally, for a given allocation $x$, a Voronoi partition $\mathcal{V}_{i}(x;\mathcal{C})$ of space $\mathcal{C}$ is defined as follows:
@@
When ties occur (i.e., when several agents are at the same minimal distance from some location), agents are prioritized lexicographically.
We assume that the locations an agent can perceive are limited to those within their Voronoi region, that is $\mathcal{X}_{i}(x)=\mathcal{V}_{i}(x;C)$. We also denote the neighborhood of $i$ as the agents connected in the dual Delaunay graph of the Voronoi partition of $i$, i.e., the neighborhood of $\mathcal{N}_{i}\subseteq\mathcal{N}$ are the agents $j\in\mathcal{N}$ whose Voronoi region is connected to the Voronoi region of $i$.
The utility function of agent $i$ is then defined as follows:
$\left.\begin{array}[]{l}u_{i}(x)=\underset{\begin{subarray}{c}c\in\mathcal{X}_{i}(x)\end{subarray}}{\sum}v_{c}\leavevmode\nobreak\ g(|x_{i}-c|).\end{array}\right.$ (11)
Note that these utility functions satisfy the assumptions of Equation (6) and Equation (7). Voronoi partition plays an important role in distributed coverage control because agents can compute the best responses improving the objective function locally, with limited communication. The process is then just a sequence of best response updates (in the sense defined above) of the different agents to compute their next locations in their local Voronoi region. Once agents are assigned these new locations, Voronoi regions are updated and the process iterates, until convergence. However, in the non-convex discrete setting the move of an agent within its partition can affect not only neighbors (as in the continuous setting), but the whole set of agents in the worst case Yun and Rus (2014). (We will see an example later in Figure 5). In theory, the guarantee of convergence requires avoiding moves that could impact beyond an agent’s neighbors Yun and Rus (2014); Sadeghi et al. (2020).
@@
(Ex. 1, cont.). Figure 2 indicates the Voronoi region of each agent by coloring the locations in the same color. (Recall that ties are broken lexicographically). Figure 2 gives the corresponding neighboring graph: agent $e$ has 3 neighbors $\{c,d,f\}$. The utilities of the agents are as follows: $u_{a}=1$, $u_{b}=1.5$, $u_{c}\simeq 3.2$ $u_{d}=4.2$ $u_{e}\simeq 5$ and $u_{f}=1.5$
@@ 3. A Neighborhood Optimal Algorithm
To address the inefficiency issue of coverage control through inter-agent communication, we extend the solution for 1-dimensional space Marden (2016) to general settings by making agents share additional information. We first present our main result on convergence and approximation ratio and then discuss the details of the algorithm.
3.1. High Level Description of the Algoritm
The idea of the algorithm is to incrementally compute an allocation $x$ based on a partition $\mathcal{P}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{N})$ of the space, which is not necessarily a Voronoi partition. Note that the neighborhood $\mathcal{N}_{i}$ is also defined over $\mathcal{P}$, instead of Voronoi partition. The algorithm is anytime and updates a solution $(x,\mathcal{P})$ for each iteration.
As in the case of Voronoi-based control, agents’ utilities $u_{i}(x,\mathcal{P})$ are localized and depend on the current partitioning of the space, i.e., we have $\mathcal{X}_{i}(x)=\mathcal{P}_{i}(x;C)$. The minimum utility enjoyed by any agent of the system is $u_{\min}(x,\mathcal{P})=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\min}\leavevmode\nobreak\ u_{i}(x,\mathcal{P})$, and the agent with this minimum utility is denoted as $i_{\min}(x,\mathcal{P})$. When there are multiple candidates for $i_{\min}$ it can be resolved using a tie-breaking rule. We first discuss the case of single $i_{\min}$. The case of multiple candidates of $i_{\min}$ is discussed later.
We follow Marden Marden (2016) and define the maximum gain in $G$ when adding $k$ new agents into space $\mathcal{C}$, as:

### 08 Apr 2024 v3

@@ Abstract.
Conventional distributed approaches to coverage control may suffer from lack of convergence and poor performance, due to the fact that agents have limited information, especially in non-convex discrete environments. To address this issue, we extend the approach of Marden (2016) which demonstrates how a limited degree of inter-agent communication can be exploited to overcome such pitfalls in one-dimensional discrete environments. The focus of this paper is on extending such results to general dimensional settings. We show that the extension is convergent and keeps the approximation ratio of 2, meaning that any stable solution is guaranteed to have a performance within 50% of the optimal one. The experimental results exhibit that our algorithm outperforms several state-of-the-art algorithms, and also that the runtime is scalable.
@@ 1. Introduction
It is known that, even in a centralized context, finding an optimal solution for the coverage problem is an NP-hard problem Megiddo and Supowit (1984). Hence, most studies focus on approximate approaches. In distributed settings, game-theoretical control approaches seek to design agents that will be incentivized to behave autonomously in a way that is well-aligned with the designer’s objective. This strategy has proven to be successful in a number of applications (see, e.g. Douchan et al. (2019)). The situation is made more difficult in practice since agents often have restricted sensing and communication capabilities. Consequently, agents must make decisions based on local information about their environment and the other agents. Unfortunately, algorithms based on local information may suffer from lack of convergence and degraded performance due to miscoordination between the agents. As for the convergence issue, it is known that a move of an agent can affect the cost of agents outside of the neighborhood, and thus, the decrease of the global cost cannot be guaranteed locally, especially in the case of discrete non-convex environment Yun and Rus (2014). The degradation of performance can be explained with a worst-case scenario in which only a single agent can perceive a large number of valuable locations within an environment, while a number of other agents cannot perceive these locations.
@@
Indeed, different settings exist depending on the assumed communication model, that is, how information can be shared within the system in order to potentially coordinate moves:
@@
    agents may not communicate any information and can only be guided by their local perception;
@@
    agents may communicate to their neighbors only (where neighbors can be defined as agents in a limited range, or connected via a network of communication);
@@
    agents may communicate beyond their neighbors, either via broadcasting, or indirectly via gossip protocols.
@@
Existing game-theoretical approaches can be classified into these types: classical Voronoi-based best-response approaches fall into either (1) or (2), depending on the assumption. Two recent approaches in type (3) explore different directions. Sadeghi et al. Sadeghi et al. (2020) proposed a distributed algorithm for non-convex discrete settings in which agents have the possibility to coordinate a move with a single other agent (meaning that an agent moves and assumes another agent takes her place simultaneously), possibly beyond their neighbors, when individual moves are not sufficient. However, the algorithm sacrifices the approximation ratio for convergence. Another related distributed approach proposed in Durham et al. (2011) is to aim for pairwise optimal partitions (which, they show, are also Voronoi partitions with the additional property that no pairs of agents can cooperate to find a better partitioning of their two regions). Again, while their approach ensures convergence (to good solutions in practice), it does not come with a guaranteed approximation ratio. In a similar spirit, Marden Marden (2016) allows coordinated moves with several (possibly, beyond two) agents, but only under the restriction that these agents are within the neighborhood. While this approach achieves a convergent algorithm with an approximation ratio of 2, by sharing only the information of the minimum utility among agents, it is limited in that the only investigated case is the one-dimensional (line) environment. This severely limits real-world applications.
Other studies rely on different techniques. Distributed constraint optimization is a natural way to model this problem Zivan et al. (2015), and several standard algorithms can be exploited. However, they do not offer approximation guarantees either. Finally, it is possible to try to achieve global optimality by developing approaches akin to simulated annealing. For example, Arslan et al. (2007) attains global optimum with Spatial Adaptive Play (SAP, a.k.a Boltzmann exploration) and uses random search to escape from the local optimum. However, this approach suffers from a slow convergence rate when the search space is large. There is also no discussion about the relationship between information and efficiency. Note that we focus on the coverage problem, and multiagent path-planning issues Galceran and Carreras (2013); Stern et al. (2019); Okumura and Défago (2022) are out of the scope of the paper.
@@
In this paper, we extend the algorithm of Marden (2016) to any-dimensional, non-convex discrete space and compare this approach with the aforementioned alternative variants of game-theoretical control. The remainder of this paper is as follows. Section 2 introduces the model and existing approaches. In Section 3 we detail the algorithm and prove that it guarantees convergence to a neighborhood optimum solution with an approximation ratio of 2 without any restriction on the dimensionality of the environment, i.e., the same guarantee as in the 1D case. Additionally, we propose a scalable extension of the algorithm and discuss the computation and communication complexity. Experiments reported in Section 4 indeed show that our algorithm outperforms existing ones, along with adhering to the theoretical approximation ratio. Furthermore, the runtime results confirm the scalability of the proposed algorithm.
@@ 2.1. Coverage Problems
We start with a set of agents $\mathcal{N}=\{1,\ldots,n\}$ and a set of resources $C=\{c_{1},\ldots,c_{m}\}$. In a coverage problem, resources are locations (or points) in a connected metric space. We assume that this is discrete finite space that is modelled as a connected graph $(C,\mathcal{E})$, where $\mathcal{E}$ is the set of edges that connect two adjacent resources. Though our approach can be extended to continuous settings, we omit the detail due to the space limit. We denote the distance between two resources $a,b\in C$ as $|a-b|$ that is the length of the shortest path.
@@
Let $g:\mathbb{R}_{+}\to\mathbb{R}_{+}$ denote a non-increasing function, $v_{c}\in\mathbb{R}_{+}$ be the weight of resource $c\in C$, and $\mathcal{C}\subseteq C$ be a partial space. Then, the objective function for $\mathcal{C}$ is defined as follows:
@@
Let us then consider the coverage problem depicted in Figure 1. For the sake of exposure, let us assume that $g(d)=1/(1+d)$, and ||$\cdot$|| ⋅ | is Manhattan distance. Agents are identified by letters $a,b,\dots f$. The environment is a grid world, where each junction is a resource (i.e. location). Circled locations are valued $v_{c}=1$, while the others are valued $v_{c}=0$. The locations covered by an agent are represented in grey and we indicate the name of the agent. For unoccupied locations, we indicate on the node the (Manhattan) distance to the closest agent. Thus, a covered location gives a utility of 1, while a location at a distance of 1 from the closest agent gives a utility 1/2, and so on. The depicted allocation $x$ has value $G(x)=1+1+\frac{1}{2}+\dots=16.4$. It is clearly sub-optimal. The reader can check that the allocation $x^{\prime}$ where agent $e$ has moved one location up would induce $G(x^{\prime})\simeq 16.8$.
@@ 2.2. Game-Theoretic Control: Generalities
This paper will focus on providing each agent $i$ with a utility function of the form $u_{i}:\mathcal{X}\to\mathbb{R}$ that will ultimately guide their individual behavior. In the rest of the paper, we formulate the set of choices of agent $i$ as $\mathcal{X}_{i}(x)$ assuming that it depends on a given allocation of all agents. For simplicity, we denote $\mathcal{X}(x)=\underset{\begin{subarray}{c}i\in\mathcal{N}\end{subarray}}{\times}\mathcal{X}_{i}(x)$.
@@
In general, the efficiency of a Nash equilibrium $G(x)$ can be smaller than the optimal value $G(x^{*})$. One of the reasons for this suboptimality is the miscoordination among self-interested agents, as agents are required to make independent decisions in response to available information. Other sources of suboptimality come from the structure of the agents’ utility functions and available choice sets. The (worst-case, among all instances) ratio between the worst Nash equilibria and the social optimum is known as the price of anarchy (PoA) Koutsoupias and Papadimitriou (1999). Here, the choice set $\mathcal{X}_{i}$ could encode physical constraints on choices, i.e. an agent can only physically choose among a subset of choices given the behavior of the collective $x$. Alternatively, the choice set could encode information availability of the agent, e.g. it is the set of choices for which the agent can evaluate its utility. Regardless of the interpretation, it is important to highlight that the structure of the choice sets can significantly alter the structure and efficiency of the resulting equilibria which we discuss in the next section.
@@ 2.3. Local Information
Several approaches in game theory seek to exploit the fact that agents typically have limited sensing power and only a local view of the situation. Marden Marden (2016) analyzed how miscoordination among agents with limited information leads to inefficient Nash equilibria. To this end, he introduced the concept of information set which is the set of choices each agent can perceive and compute the resulting utilities. In this paper, $\mathcal{X}_{i}(x)$ corresponds to the information set that is the set of locations agent $i$ can perceive based on spatial proximity. With this notion, an allocation is a Nash equilibrium (Equation  3) if agents are at least as happy as any choice for which they can evaluate their utility. Observe that the information set $\mathcal{X}_{i}(x)$ is a state-dependent notion: the local information available may vary depending on the current allocation $x$.
@@ 2.4. Linking Information and Inefficiency
Intuitively, local information can cause distributed systems based on game-theoretic control to get stuck in an inefficient allocation. Indeed, some agents who may obtain higher utilities for a resource may not have access to this information. The redundancy index hence gives insights into the amount of information available to the agents and about guarantees on the worst-case ratio.
Marden Marden (2016) investigated this interplay, in the broader context of resource allocation games. The global objective function $G$ is assumed to be monotone submodular i.e., it satisfies the following conditions:
@@
The following theorem shows how the value of $f$ impacts the efficiency of Nash equilibrium allocations:
@@
Thus, we know that the approximation ratio of a distributed allocation algorithm can be 2 in the case of full information ($f=n$), because $G(x)\geq\frac{1}{2}G(x^{*})$ in this case (from Equation 8). Also, it is impossible to guarantee an approximation ratio better than $n$ in case of $f=1$ (from Equation 9). We will use this result in the later section to analyze the efficiency of coverage control algorithms.
@@ 2.5. Voronoi-Based Control
Most standard approaches for coverage control are based on Voronoi partitioning Gao et al. (2008). A Voronoi partition divides the space into local regions for each agent. Formally, for a given allocation $x$, a Voronoi partition $\mathcal{V}_{i}(x;\mathcal{C})$ of space $\mathcal{C}$ is defined as follows:
@@
When ties occur (i.e., when several agents are at the same minimal distance from some location), agents are prioritized lexicographically. Let $\mathcal{P}=(\mathcal{P}_{1},\ldots,\mathcal{P}_{N})$ denote a partition of the space. In the rest of the paper, we define the utility function of agent $i$ depending on partition $\mathcal{P}$ as follows:
$\left.\begin{array}[]{l}u_{i}(x,\mathcal{P})=\underset{\begin{subarray}{c}c\in\mathcal{P}_{i}\end{subarray}}{\sum}v_{c}\leavevmode\nobreak\ g(|x_{i}-c|).\end{array}\right.$ (11)
In this subsection, we assume that the locations an agent can perceive are limited to those within their Voronoi region, that is $\mathcal{P}_{i}=\mathcal{X}_{i}(x)=\mathcal{V}_{i}(x;C)$. Therefore, the utility function (Equation 11) satisfies the assumptions of Equation 6 and Equation 7. We also denote the neighborhood of $i$ as the agents connected in the dual Delaunay graph of the Voronoi partition of $i$, i.e., the neighborhood of $\mathcal{N}_{i}\subseteq\mathcal{N}$ are the agents $j\in\mathcal{N}$ whose Voronoi region is connected to the Voronoi region of $i$. Voronoi partition plays an important role in distributed coverage control because agents can compute the best responses improving the objective function locally, with limited communication. The process is just a sequence of best response updates (in the sense defined above) of the different agents to compute their next locations in their local Voronoi region. Once agents are assigned these new locations, Voronoi regions are updated and the process iterates, until convergence. However, in the non-convex discrete setting the move of an agent within its partition can affect not only neighbors, but the whole set of agents in the worst case Yun and Rus (2014). We will see an example later in Figure 5. In theory, the guarantee of convergence requires avoiding moves that could impact beyond an agent’s neighbors Yun and Rus (2014); Sadeghi et al. (2020).
@@
(Ex. 1, cont.). Figure 2 indicates the Voronoi region of each agent by coloring the locations in the same color. Recall that ties are broken lexicographically. Figure 2 gives the corresponding neighboring graph: agent $e$ has 3 neighbors $\{c,d,f\}$. The utilities of the agents are as follows: $u_{a}=1$, $u_{b}=1.5$, $u_{c}\simeq 3.2$ $u_{d}=4.2$ $u_{e}\simeq 5$ and $u_{f}=1.5$
As mentioned in the introduction, the communication requirements of these approaches is very low, as agents never need to exchange beyond their neighbors in the Delaunay graph (note that this can also be achieved with a range-limited gossip protocol, as long as an appropriate motion protocol is designed Durham et al. (2011)). In the next section, we introduce our protocol which assumes agents can always communicate with their neighbors, and requires them in addition to build a communication tree to spread a limited amount of information to facilitate coordination. Going beyond this communication model is left for future work.
@@ 3. A Neighborhood Optimal Algorithm
To address the inefficiency issue of coverage control through inter-agent communication, we extend the solution for 1-dimensional space Marden (2016) to general settings by making agents share additional information.
3.1. Key Notions
The idea of the algorithm is to incrementally compute an allocation $x$ based on a partition $\mathcal{P}$, which is not necessarily a Voronoi partition. Note that the neighborhood $\mathcal{N}_{i}(\mathcal{P})$ is also defined over $\mathcal{P}$, as the set of agents that have partitions adjacent to $\mathcal{P}_{i}$. The algorithm is anytime and updates a solution $(x,\mathcal{P})$ for each iteration.
We follow Marden (2016) and define the maximum gain in $G$ when adding $k$ new agents into space $\mathcal{C}$, as: